# Prediction of flow and elastic stresses in a viscoelastic turbulent channel flow using convolutional neural networks

### v1

@@ 4 Conclusions
The present work highlights the capability of a data-driven approach to perform non-intrusive sensing in viscoelastic turbulent flows. Here we demonstrate the ability of CNN-based models to accurately reconstruct the velocity fluctuations in viscoelastic turbulence close the wall, utilizing the two wall-shear fluctuation components and the wall-pressure fluctuations as inputs. Additionally, the network models successfully reproduce the polymeric-stress fluctuation fields from the DNS velocity-fluctuation fields. Moreover, the feasibility of these network models to extract polymer stress fluctuation fields of interest solely from wall input fluctuations and predicted velocity-fluctuations is explored. Overall, the network effectively reconstructs the large-scale features of the polymer-stress fields using wall inputs and predicted velocity fields. Furthermore, the developed models exhibit enhanced accuracy in predicting quantities of interest during the hibernation intervals, facilitating a deeper understanding of the underlying physics during low-drag events when the model is deployed in a practical application. These non-intrusive-sensing models hold valuable applications in experimental settings (Vinuesa et al., 2023), enabling the construction of polymeric stresses in turbulent flows from velocity fields or wall-inputs, which otherwise would be challenging or impossible to quantify experimentally.

### v2

@@ 4 Conclusions
The present work highlights the capability of a data-driven approach to perform non-intrusive sensing in viscoelastic turbulent flows. Here we demonstrate the ability of CNN-based models to accurately reconstruct the velocity fluctuations in viscoelastic turbulence close to the wall, utilizing the two wall-shear fluctuation components and the wall-pressure fluctuations as inputs. Additionally, the network models successfully reproduce the polymeric-stress fluctuation fields from the DNS velocity-fluctuation fields. Moreover, the feasibility of these network models to extract polymer-stress-fluctuation fields of interest solely from wall input fluctuations and predicted velocity fluctuations is explored. Overall, the network effectively reconstructs the large-scale features of the polymer-stress fields using wall inputs and predicted velocity fields. Furthermore, the developed models exhibit enhanced accuracy in predicting quantities of interest during the hibernation intervals, facilitating a deeper understanding of the underlying physics during low-drag events when the model is deployed in a practical application. These non-intrusive-sensing models hold valuable applications in experimental settings (Vinuesa et al., 2023), enabling the determination of polymeric stresses in turbulent flows from velocity fields or wall inputs, which otherwise would be challenging or impossible to quantify experimentally. However, accurately predicting the polymer stresses from wall inputs is found to be more challenging than predicting velocity fields from the same, since wall inputs typically do not contain fingerprints of the smallest scales, and we show that small scales in velocity fields are connected to a wide range of scales in polymeric stress fields.
Aimed towards extracting polymer-stress information from velocity-fluctuation fields in a possible experimental investigation of viscoelastic turbulent channel flow, a number of FCN models are trained with inputs corresponding to different thresholds of small-scale wavelengths in the velocity-fluctuation fields. We find that accurately capturing the turbulent kinetic energy in the near-wall fields is crucial for retrieving second-order turbulent statistics of polymer stresses. Specifically, experimental acquisition of velocity fluctuations would in principle need to resolve the finer scales, smaller than 10 viscous units, to obtain more reliable polymer stress behaviour at the smallest scales, which determines the accurate rate of energy transfer from flow to polymers. We show that this in turn facilitates an accurate estimation of the RMS of polymer stresses in the buffer region of viscoelastic turbulent channel flow.
Concluding, the results demonstrate the potential of data-driven models to predict instantaneous fields in non-Newtonian wall turbulence, which is useful for flow control or estimation of stress fluctuations that cannot be measured. As a next step, de-noised experimentally measured velocity fields and wall pressure measurements could be given as inputs to the network that has been trained by simulations. Furthermore, direct numerical simulations with filtered velocity fluctuations can be performed to further isolate the physical mechanisms behind how small-scale velocity fluctuations influence the polymer stress fluctuation fields at a wide range of scales.
Appendix A Performance metrics for network models
In the present study, hyperparameter tuning was conducted to investigate the sensitivity of various parameters, including network model capacity, depth, kernel size, batch size, and learning rate, on the errors in predicting the RMS of velocity and polymer stress fluctuations. Approximately 15 different training runs were performed to optimize these parameters, and a more thorough exploration of the hyperparameter space could further enhance prediction performance. In the present study, a fully convolutional neural network containing kernels and 31 hidden layers constituting trainable parameters is employed (refer figure 1). The initial weights for the network are distributed randomly with a Gaussian distribution and a scheduled learning rate () was provided to the Adam algorithm (Kingma & Ba, 2015) in the form , with corresponding to initial learning rate and are the tunable parameters denoted by learning rate drop and epoch drop, respectively (here, epoch corresponds to one entire pass of training data). In this study, , and and a batch size of 16 was employed in the training of FCN.
Additionally, U-Net (with five skip connections) and Generative Adversarial Network (GAN) models (with about 2 million trainable parameters in the generator) were also trained, and corresponding evaluations revealed comparable performance in terms of accuracy to that achieved with the FCN in the present study. Hence, the FCN with the optimal training parameters as obtained with the exploratory study is employed in the present work.
The prediction errors for V-predictions are provided in table 1 and various metrics for E-predictions and V-E-predictions are detailed in table 2. The error metrics reported in table 1 and table 2 are obtained from the samples contained in the test dataset and averaged over three different training runs for the FCN model. The variances are indicated in the braces.
Parameters | ||||
15 | 0.14 ( 0.01) | 0.32 ( 0.11) | 0.41 ( 0.02) | |
30 | 0.29 ( 0.01) | 0.44 ( 0.06) | 0.53 ( 0.02) | |
50 | 0.47 ( 0.01) | 0.79 ( 0.15) | 0.59 ( 0.04) | |
15 | 2.62 ( 1.2) | 13.11 ( 2.7) | 8.02 ( 3.8) | |
30 | 5.21 ( 0.3) | 15.90 ( 4.5) | 10.62 ( 1.9) | |
50 | 12.73 ( 3.9) | 24.72 ( 6.5) | 14.72 ( 4.4) | |
15 | 0.996 ( 0.003) | 0.889 ( 0.040) | 0.932 ( 0.018) | |
30 | 0.942 ( 0.002) | 0.763 ( 0.015) | 0.771 ( 0.010) | |
50 | 0.811 ( 0.001) | 0.623 ( 0.011) | 0.643 ( 0.010) |
Parameters | |||||
15 | 0.19 ( 0.01) | 0.20 ( 0.01) | 0.45 ( 0.01) | 0.63 ( 0.01) | |
30 | 0.21 ( 0.01) | 0.20 ( 0.02) | 0.49 ( 0.02) | 0.60 ( 0.01) | |
50 | 0.19 ( 0.01) | 0.17 ( 0.02) | 0.48 ( 0.01) | 0.55 ( 0.01) | |
15 | 6.03 ( 1.8) | 8.68 ( 1.4) | 42.52 ( 1.0) | 54.71 ( 2.6) | |
30 | 11.54 ( 2.7) | 9.47 ( 1.8) | 49.40 ( 3.4) | 51.14 ( 0.7) | |
50 | 11.46 ( 1.4) | 5.52 ( 1.4) | 62.37 ( 3.2) | 49.54 ( 4.8) | |
15 | 0.906 ( 0.004) | 0.924 ( 0.007) | 0.484 ( 0.006) | 0.362 ( 0.013) | |
30 | 0.906 ( 0.006) | 0.917 ( 0.010) | 0.296 ( 0.013) | 0.433 ( 0.008) | |
50 | 0.905 ( 0.007) | 0.924 ( 0.011) | 0.207 ( 0.012) | 0.426 ( 0.015) |
Appendix B Effects of low-pass filtering the velocity fluctuations on prediction of polymer shear stress
A sample instantaneous field of the polymer shear stress () in the test dataset is shown in figure 15. The sampled DNS field at different wall-normal locations serves as a reference while the corresponding predictions obtained from FCN with respective inputs at different cut-off wavelengths are depicted. The figure shows that the resulting predictions from FCN lack certain small-scale features with increasing for different wall-normal locations, a fact that is attributed to the absence of fine-scale features in the input fields.
Further, the distribution of energy in different scales for both the DNS and the FCN-predicted polymer-shear stress field at is illustrated in figure 15, which also identifies that the resulting predictions from FCN increasingly feature energy-containing large-scale structures and lack energy content in small scales with increasing .
Appendix C Joint probability density function of polymer shear stress
A similar observation can be made for the prediction of the polymer shear stress with respect to turbulent shear stress as outlined in §3.6 for the trace of polymer stress. The probability density function for the RMS-normalized polymer shear stress is shown in figure 16. The predictions from FCN using DNS input velocity-fluctuations closely capture the probability density function observed with the DNS polymer shear stress fields in the test dataset. However, utilizing the filtered velocity-fluctuation fields with results in the distribution of the data to be closer to the mean value, indicating an under-prediction of the fluctuations of the polymer shear stress.
From figure 16, observing the distribution from DNS samples, we find the polymer shear stress to mitigate the production of turbulent shear stress. Furthermore, predictions of polymer shear stress using filtered fields are under-predicted, especially in regions with negligible turbulent shear stress. Filtering the input fields alters the Reynolds shear stress, leading to increased errors in the FCN’s predictions of polymer shear stress.